HUBBLE SPACE TELESCOPE DAILY REPORT #2266 PERIOD COVERED: 0000Z (UTC) 12/08/98 - 0000Z (UTC) 12/09/98 Daily Status Report as of 343/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Ten Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 8081 (STIS CCD Hot Pixel Annealing -- II) The Space Telescope Imaging Spectrograph (CCD) was used to make a dark and a bias calibration observation. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD/MA1/MA2 7400 (Evidence from Boron for Massive Star Evolution) The Space Telescope Imaging Telescope (CCD/MA1/MA2) was used to observe the B III and B II resonance lines at 2065.8 and 1362.5 Angstroms, respectively, in the slightly evolved blue star PG0832+676 using the HST STIS MAMA detectors and echelle gratings. The proposal completed with no reported problems. 1.5 Completed Two Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.6 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.7 Completed Two Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.8 Completed STIS/MA1 7298 (A Snapshot Survey for UV bright Quasars at redshift below 3) The Space Telescope Imaging Spectrograph (MA1) was used to make ultraviolet snapshot observations of the bright quasar PKS0858-27 which has a redshift (Z) below 3.0. The goal is to find quasars bright enough in the ultraviolet for spectroscopic follow-up in the range of 900-1200 Angstroms. A wave length calibration image was also taken. The observation were completed as scheduled, and no anomalies were reported. 1.9 Completed STIS/MA2 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 6344 (Low - Mass Binaries in the Hyades - Completing the Survey) The WF/PC-2 was used to make observations of the low mass binary star RHY164 (M-Class Dwarf Stars) which is located in the Hyades Cluster. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Two Sets of WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.12 Completed WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected AGNs) The WF/PC-2 was used to make high resolution observations of the known low red shifted x-ray active galactic nuclei (AGN) MS1233.3+7. These data will allow a study of the optical properties of an unbiased sample of AGN host galaxies. The observations were completed as scheduled, and no problems were reported. 1.13 Completed NIC/1/2/3 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera One, Two and Three) was used to make a corographic search for massive planets around the nearby, young main sequence star TWA6B. Because of the extreme youth of this object, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. Two TTRs were written for a total of five IP data hits on days 342 and 343. Another TTR documented four data inversions for day 342. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH